If f(x)=2x+1 and g(x)=x-12, then f(g(x)) =??

Thank you for your help!

Answer by nerdybill(7384)

If f(x)=2x+1 and g(x)=x-12, then f(g(x)) =??
.
Since:
g(x)=x-12
then,
we substitute the 'x' from f(x) with "x-12" thus:
f(g(x)) = 2(x-12)+1
f(g(x)) = 2x-24+1
f(g(x)) = 2x-23 (this is what they're looking for)